The government wants to test out their new rockets. The path of their rockets are given (measured in kilometers) by h = -0.5d(2d - 500,000). h represents height, and d represents horizontal distance.
a) What is the horizontal distance that the rockets travels?
b) What is the maximum height the rocket flies?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The government wants to test out their new rockets. The path of their rockets are given (measured in kilometers) by h = -0.5d(2d - 500,000). h represents height, and d represents horizontal distance.
a) What is the horizontal distance that the rockets travels?
Assuming the launch and impact are the same height:
h = -d^2 + 250000d = 0
d = 250000 kms - not realistic
====================
b) What is the maximum height the rocket flies?
Max height is the "half-way" point of horizontal distance
h = -(125000)^2 + 250000*125000 = 15625000000 kms - even less realistic

Answer by ikleyn(52780)

.

It looks like this government believes that the Earth is flat and has the diameter at least 250,000 kilometers.


Thanks for making me laugh.